# An Image is Worth More Than 16$\times$16 Patches: Exploring Transformers on Individual Pixels

### v1

@@ 5 Experiments for PiT
PiT-S 86.4 96.6
ViT-B/2 [58] 72.6 -
(a) CIFAR-100 classification.
model Acc@1 Acc@5
@@
PiT-B 76.1 92.6
(b) ImageNet classification.
Table 3: Results for case study #1 (supervised learning). Learning from scratch, we compare PiT and ViT [22] (with patch size 2$\times$2). We report results on (a) CIFAR-100 [38]: with 32$\times$32 inputs, PiT substantially outperforms ViT. Note that our ViT baselines are already well-optimized, e.g. [58] reports 72.6% when training from scratch with a larger model; (b) ImageNet [19]: using the training pipeline [61, 26] highly optimized for ViT, we still observe accuracy gains with PiT. Due to computation constraints, we use an input size of 28$\times$28 (lower than [22, 61]).
@@ 5.1 Case Study #1: Supervised Learning
In this study, we train and evaluate PiT from scratch without any pre-training [26, 10]. Our baselines are ViTs with patch sizes 2$\times$2.
@@
We use two datasets: CIFAR-100 [38] with 100 classes and 60K images combined and ImageNet [19] with 1K classes and 1.28M images for training, 50K for evaluation. While CIFAR-100 is suitable for exploring the effectiveness of PiT due to its intrinsic image size of $32\times 32$, ImageNet has many more images which helps us to further confirm our findings.
@@
For both datasets, we train our models on the train split and report the top-1 (Acc@1) and top-5 (Acc@5) accuracy on the val split.
Implementation details.
For CIFAR-100, due to the lack of optimal settings even for ViT, we search for the recipe and report results using model sizes Tiny and Small. We use the augmentations from the released demo of [27] to train from scratch, as we found more advanced augmentations (e.g., AutoAug [13]) not helpful in this case. All models are trained using AdamW [45] with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$. We use a batch size of 1024, weight decay of 0.3, drop path [33] of 0.1, and initial learning rate of 0.004 which we found to be the best for all the models. We use a linear learning rate warm-up of 20 epochs and cosine learning rate decay to a minimum of 1e-6. Our training lasts for 2400 epochs, compensating for the small size of dataset.
On ImageNet, we closely follow the from-scratch training recipe from [61, 26] for ViT and report results using PiT-S and PiT-B. Due to the limit in computation, images are crop-and-resized to $28{\times}28$ as the low-resolution inputs by default. Globally average pooled outputs are used for classification. The training batch size is 4096, initial learning rate is $1.6{\times}10^{-3}$, weight decay is 0.3, drop path is 0.1, and training length is 300 epochs. MixUp [65] (0.8), CutMix [64] (1.0), RandAug [14] (9, 0.5), and exponential moving average (0.9999) are used.
Main comparisons (Tab. 3).
While our baselines for both ViT variants (ViT-T and ViT-S) are well-optimized on CIFAR-100 (e.g., [58] reports 72.6% Acc@1 when training from scratch with ViT-B, whilst we achieve 80+with smaller sized models), PiT-T improves over ViT-T by 1.5% of Acc@1; and when moving to the bigger model (S), PiT shows an improvement of 1.3% of Acc@1 over the small model (T) while ViT seems to be saturated. These results suggest compared to the patch-based ViT, PiT is potentially learning new, data-driven patterns directly from pixels.
@@
Figure 2: Two trends for PiT vs. ViT on ImageNet. Since PiT can be viewed as ViT with patch size 1$\times$1, the trends w.r.t. patch size is crucial to our finding. In (a), we vary the ViT-B patch size but keep the sequence length fixed (last data point equivalent to PiT) – so the input size is also varied. While Acc@1 remains constant in the beginning, the size, or the amount of information quickly becomes the dominating factor that contributes to the degeneration, and PiT is the worst. On the other hand, in (b) we vary the ViT-S patch size while keeping the input size fixed. The trend is opposite – reducing the patch size is always helpful and the last point (PiT) becomes the best. The juxtaposition of these two trends gives a more complete picture of the relationship between input size, patch size and sequence length.
Our observation also transfers to ImageNet – albeit with a significantly lower resolution our results are significantly lower than the state-of-the-art [22, 61] (80+%), PiT still outperforms ViT in both settings we have experimented.
PiT vs. ViT: a tale of two trends.
@@
If position embeddings are learned, PiT is simply a version of ViT with 1$\times$1 patches. Therefore, it is crucial to study the performance trend when varying the patch sizes in ViT. There are three variables in concern: sequence length ($L$), input size ($H{\times}W$) and patch size ($p$). They have a deterministic relationship: $L=H{\times}W/(p^{2})$. Thus we have two ways to study the Acc@1 trend w.r.t. patch size $p$:
@@
  • Fixed sequence length. We show the trend on ImageNet with a fixed $L$ in Fig. 2(a). The model size is ViT-B. In this plot, the input size varies (from 224$\times$224 to 14$\times$14) as we vary the patch size (from 16$\times$16 to 1$\times$1). The last data point is equivalent to PiT-B. If we follow this trend, then PiT is the worst. This means sequence length is not the only deciding factor for Acc@1 even for classification – a task where a single label is assigned to the entire image. Input size, or the amount of information fed into the model is arguably a more important factor, especially when the size is small. It’s only when the input size is sufficiently large (e.g., 112$\times$112), the additional benefit of further enlarging the size starts to diminish. This also means when the amount of information from the input is not enough, PiT– or any architecture that follows this design (e.g., iGPT [7]) would not work well.
@@
  • Fixed input size. Our finding resides in the other trend, when we fix the input size (therefore the amount of information), and vary the patch size on ImageNet in Fig. 2(b). The model size is ViT-S. Interestingly, we observe an opposite trend here: it is always helpful to decrease the patch size (or increase the sequence length), aligned with the prior studies that claim sequence length is highly important. Note that the trend holds even when it ultimately reaches PiT – a model without any design for locality. So PiT performs the best in accuracy compared to ViTs.
With these two trend figures in Fig. 2, our study augments the observations made from previous studies, as they mainly focused on regimes where the input size is sufficiently large [2, 32], and presents a more complete picture.
To see what PiT has learned, we show visualizations of the attention maps, position embeddings from PiT in Appendix 0.A.
model pre-train Acc@1 Acc@5
PiT-T 85.1 96.4
@@
Table 4: Results for case study #2 (self-supervised learning). We use PiT and ViT [22] (patch size 2$\times$2) for MAE pre-training [26] on CIFAR-100, and then fine-tune with supervision. For both (a) Tiny and (b) Small sized variants, pre-training boosts performance. PiT-S offers a bigger gap over ViT-S, suggesting PiT can scale better.
5.2 Case Study #2: Self-Supervised Learning
In this subsection, we study PiT with self-supervised pre-training and then fine-tuning for supervised classification. In particular, we choose MAE [26] due to its efficiency that only retains 25% of the sequence length for the encoder, and its effectiveness for fine-tuning based evaluation protocols.
Datasets.
We use CIFAR-100 [38] due to its inherent size of $32\times 32$ for images. This allows us to fully explore the use of pixels as tokens on the original resolution.
@@ 5.2 Case Study #2: Self-Supervised Learning
We first perform pre-training on the train split. Then it serves as the initialization in the fine-tuning stage (also trained on train). Again, we use image classification on CIFAR-100 as the downstream task and report the top-1 (Acc@1) and top-5 (Acc@5) accuracy on the val split.
Implementation details.
We follow standard MAE and use a mask ratio of 75% and select tokens randomly. Given the remaining 25% visible tokens, the model needs to reconstruct masked regions using pixel regression. Since there is no known default setting for MAE on CIFAR-100 (even for ViT), we search for recipes and report results using PiT-T and PiT-S. The same augmentations as in [26] are applied to the images during the pre-training for simplicity. All models are pre-trained using AdamW with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$. We follow all of the hyper-parameters in [26] for the pre-training of 1600 epochs except for the initial learning rate of 0.004 and a learning rate decay of 0.85 [11]. Thanks to MAE pre-training, we can fine-tune our model with a higher learning rate of 0.024. We also set weight decay to 0.02, layer-wise rate decay to 0.65, and drop path to 0.3, $\beta_{2}$ to 0.999, and fine-tune for 800 epochs. Other hyper-parameters closely follow the scratch training recipe for supervised learning (see Sec. 5.1).
Figure 3: Qualitative results for case study #3 (image generation). These 256$\times$256 samples are from PiT-L trained on ImageNet, following the same architecture design and generation protocol as that of DiT [51]. [^3] The models were not prompted to generate using any of the people images/classes from ImageNet (scuba diver, baseball player, bridegroom). PiT generations include fine features and of reasonably similar quality as the DiT generations with locality [51].
Results.
As shown in Tab. 4, we find that for PiT too, self-supervised pre-training with MAE improves accuracy compared to training from scratch. This is true for both PiT-T and PiT-S that we experimented with. Notably, the gap between ViT and PiT, with pre-training, gets larger when we move from Tiny to Small models. This suggests PiT can potentially scale better than ViT.
5.3 Case Study #3: Image Generation
We switch to image generation with Diffusion Transformer (DiT) [51] in this section, which has a modulation-based architecture different from vanilla ViT, and operates on the latent token space from VQGAN [24, 54] that shrinks the input size by 8$\times$. Dataset-wise, we use ImageNet for class-conditional generation, and each image is center-cropped to 256$\times$256, resulting in an input feature map size of 32$\times$32$\times$4 (4 is channel dimension). PiT-L is fed with this feature map, same as its baseline DiT-L/2 [51].
Evaluation metrics.
The generation quality is measured by standard metrics: Fréchet Inception Distance (FID) [29] with 50K samples, sFID [48], Inception Score (IS) [55], and precision/recall [40], using reference batches from the original TensorFlow evaluation suite of [21].
Implementation details.
We followed the settings for DiT training, with a larger batch size (2048) to make the training faster (the original recipe uses a batch size of 256). To make the training stable, we perform linear learning rate warm up [25] for 100 epochs and then keep it constant for a total of 400 epochs. We use a maximum learning rate of 8e-4, with no weight decay applied.
@@ 5.3 Case Study #3: Image Generation
Sampled generations from PiT-L are shown in Fig. 3. The sampling takes 250 time steps, with the latent diffusion outputs mapped back to the pixel space using the VQGAN decoder. A classifier-free guidance [30] scale of 4.0 is used. All generations are detailed and reasonable compared to the DiT models with the locality inductive bias [51].
@@
DiT-XL/2 [51], no guidance 10.67 - - - -
Table 5: Results for case study #3 (image generation). We use reference batches from the original evaluation suite of [21], and report 5 metrics comparing locality-biased DiT-L/2 and PiT-L. With the exception of the last two rows, we use classifier-free guidance [30] (scale 1.5) during the generation process (250 steps). Last row is from [51], compared to which our baseline is significantly stronger (8.90 FID [29] with DiT-L, vs. 10.67 with DiT-XL and longer training). Overall, our finding generalizes well to this new task with a different architecture and different input representations.
@@
We summarize qualitative comparisons between DiT-L/2 and PiT-L in Tab. 5. First, our baseline is strong despite the change of training recipe: compared to the reference 10.67 FID [51] with a larger model (DiT-XL/2) and longer training ($\sim$470 epochs), our DiT-L/2 achieves 8.90 without classifier-free guidance. Our main comparison (first two rows) uses a classifier-free guidance of 1.5 with 250 sampling steps. With PiT operating on the latent ‘pixels’, it outperforms the baseline on three metrics (FID, sFID and IS), and is on-par on precision/recall. With extended training, the gap is bigger (see Appendix 0.B).
Our demonstration on the image generation task is an important extension of PiT. Compared to the case studies on discriminative benchmarks from Sec. 5.1 and Sec. 5.2, the task has changed; the model architecture is changed from standard ViT to a conditioned one; the input space is also changed from raw pixels to latent encodings from the VQGAN tokenizer. The fact that PiT works out-of-box suggests our observation generalizes well, and locality-free architecture can be used across different tasks, architectures, and operating representations.
@@
Figure 4: Pixel permutation to study the effect of patchification in ViT. We swap pixels within a Hamming distance of $\delta$ and do this $T$ times ($\delta=\inf$ means no distance constraint). Illustrated is a 8$\times$8 image divided into 2$\times$2 patches. Here we show permutation with $T=4$ pixel swaps (4 double-headed arrows).
6 Locality Designs in ViT
Experimental setup.
We use ViT-B for ImageNet supervised classification. We adopt the exact same hyper-parameters, augmentations, and other training details from the scratch training recipe from [26]. Notably, images are crop-and-resized to 224$\times$224 and divided into 16$\times$16 non-overlapping patches.
Position embedding.
Similar to the investigation in [10], we choose from three candidates: sin-cos [62], learned, and none. The first option introduces locality into the model, while the other two do not. The results are summarized below:
PE sin-cos learned none
Acc@1 82.7 82.8 81.2
Our conclusion is similar to the one drawn by [10] for self-supervised representation evaluation: learnable position embeddings are on-par with fixed sin-cos ones. Intrestingly, we observe only a minor drop in performance even if there is no position embedding at all – ‘none’ is only worse by 1.5% compared to sin-cos. Note that without position embedding, the classification model is fully permutation invariant w.r.t. patches, though not w.r.t. pixels – will show evidence next.
@@ 6 Locality Designs in ViT
Next, we use learnable position embeddings and study patchification. To systematically reduce locality from patchification, our key insight is that neighboring pixels should no longer be tied in the same patch. To this end, we perform a pixel-wise permutation before diving the resulting sequence into separate tokens. Each token contains 256 pixels, same in number to pixels in a 16$\times$16 patch. The permutation is shared, i.e., stays the same for all the images – including the ones for testing.
The permutation is performed in $T$ steps, each step will swap a pixel pair within a distance threshold $\delta\in[2,\inf]$ ($2$ means within the 2$\times$2 neighborhood, $\inf$ means any pixel pair can be swapped). We use hamming distance on the 2D image grid. $T$ and $\delta$ control how ‘corrputed’ an image is – larger $T$ or $\delta$ indicates more damage to the local neighborhood and thus more locality bias is taken away. Fig. 4 illustrates four such permutations.

### v2

@@ 4 Experiments
ViT-S/1 86.4 96.6
ViT-B/2 (Shen et al., 2023) 72.6 -
(a) CIFAR-100 classification
Acc@1 Acc@5
@@
ViT-B/1 76.1 92.6
ViT-L/2 75.6 92.3
ViT-L/1 76.9 93.0
(b) ImageNet classification
Acc@1 Acc@5
ViT-S/2 45.8 68.3
ViT-S/1 46.3 68.9
(c) Oxford-102-Flower fine-grained classification
RMSE ($\downarrow$) RAE ($\downarrow$)
ViT-S/2 0.80 0.78
ViT-S/1 0.72 0.74
(d) NYU-v2 depth estimation (regression)
Table 3: Results for case study #1: supervised learning. We use ViT (Dosovitskiy et al., 2021), either with non-degenerated patch size ($\times2$), or with our locality-free modification that treats each pixel as a token (in bold). We report results on (a) CIFAR-100 (Krizhevsky, 2009): the pixel variant outperforms ViT with $\times2$ patches of the same model size. Note that our baselines are already highly-optimized, e.g. Shen et al. (2023) reports significantly worse results despite larger model size; the observation generalizes well to the larger (b) ImageNet (Deng et al., 2009) dataset, fine-grained classification on (c) Oxford-102-Flower (Nilsback & Zisserman, 2008), and depth estimation (regression) on (d) NYU-v2 (Silberman et al., 2012) which further requires spatial understanding. All these suggest Transformers can not just learn, but learn well without any inductive bias of locality.
@@ 4.1 Case Study #1: Supervised Learning
In this study, we train and evaluate ViTs (Dosovitskiy et al., 2021) with or without locality for supervised learning tasks. Our baselines are ViTs with patch sizes $\times2$. For more implementation details, please see Appendix A.
@@
In total we use four datasets, with two for object classification: CIFAR-100 (Krizhevsky, 2009) has 100 classes and 60K images combined, and ImageNet (Deng et al., 2009) has 1K classes and 1.28M images for training, 50K for evaluation. CIFAR-100 is well-suited for exploring the effectiveness of locality-free architectures due to its intrinsic image size of $32\times 32$. ImageNet has many more images and is more frequently used for modern architecture design. We also validate our finding on: fine-grained classification using the Oxford-102-Flower dataset (Nilsback & Zisserman, 2008) and depth estimation using the NYU-v2 dataset (Silberman et al., 2012).
@@
For the classification datasets, we train our models on the train split and report the top-1 (Acc@1) and top-5 (Acc@5) accuracy on the val split. The root mean squared error (RMSE) and relative absolute error (RAE) are reported on the val split for depth estimation.
Main results.
As shown in Tab. 3, while our baselines for both ViT variants (ViT-T and ViT-S) are well-optimized on CIFAR-100 (e.g., Shen et al. (2023) reports 72.6% Acc@1 when training from scratch with ViT-B, whilst we achieve 80+% with smaller sized models), our locality-free variant, ViT-T/1, improves over ViT-T/2 by 1.5% of Acc@1; and when moving to the bigger model (S), ViT-S/1 leads to an improvement of 1.3% in Acc@1 over ViT-T/1, while ViT/2 starts to saturate. These results suggest compared to the patch-based ViT, ViT with our locality-free design is potentially learning new, data-driven patterns directly from individual pixels.
@@
Figure 2: Two trends with ViT. Since our Transformer on pixels can be viewed as ViT with patch size $\times1$, the trends w.r.t. patch size is crucial to our finding. In (a), we vary the ViT-B patch size but keep the sequence length fixed (last data point is locality-free) – so the input size is also varied. While Acc@1 remains constant in the beginning, the input size, or the amount of information quickly becomes the dominating factor that deteriorates accuracy. On the other hand, in (b) we vary the ViT-S patch size while keeping the input size fixed. The trend is opposite – reducing the patch size is always helpful and the last point (locality-free) becomes the best. The juxtaposition of these two trends gives a more complete picture of the relationship between input size, patch size and sequence length.
Our observation also transfers to ImageNet – though with a much lower resolution our results are also lower than the state-of-the-art (Dosovitskiy et al., 2021; Touvron et al., 2021) (80+%), the pixel-based ViT still outperforms patch-based ViT with all model sizes we have experimented, and interestingly synergizes better with larger-sized models like ViT-L. To see what the models have learned, we visualize the attention maps, position embeddings in Appendix B.
At the expense of longer sequence length, our experiments on fine-grained classification and depth estimation further confirm the effectiveness of the locality-free architecture in boosting accuracies/reducing errors. Notably, fine-grained classification requires nuanced understanding in details; whereas depth estimation demands better capture of spatial relationships. It’s impressive that Transformers can learn useful patterns purely from data, without assuming inputs are on 2D image grids.
Two trends with ViT.
@@
With learned position embeddings, our finding is simply based on a ViT variant with $\times1$ ‘patches’. Then why this is not discovered earlier? To us, the major reason is that there are two different trends with the three variables in concern: sequence length ($L$), input size ($H{\times}W$) and patch size ($p$). They have a deterministic relationship: $L=H{\times}W/(p^{2})$, and can be studied on ImageNet either via fixed sequence length or fixed input size:
@@
  • Fixed sequence length. We show this trend with a fixed $L$ in Fig. 2(a). The model size is ViT-B. In this plot, the input size varies (from $\times224$ to $\times14$) as we vary the patch size (from $\times16$ to $\times1$). If we follow this trend, then a locality-free architecture (right most) is the worst. This means sequence length is not the only deciding factor – input size, or the amount of information fed into the model is arguably a more important factor, especially when it’s small. It’s only when the size is sufficiently large (e.g., $\times112$), the additional benefit of further enlarging the size starts to diminish. This also gives an intuitive explanation that when there is no enough input information, advanced architecture or pre-training (e.g., iGPT (Chen et al., 2020a)) would struggle to recover this hard gap.
@@
  • Fixed input size. Our finding is made when following the other trend – fixing the input size to $\times64$ (thus the amount of information), and varying the patch size on ImageNet in Fig. 2(b). The model size is ViT-S. Interestingly, we observe an opposite trend here: it is always helpful to decrease the patch size (or increase the sequence length), aligned with the prior studies that find sequence length is highly important (Beyer et al., 2022; Hu et al., 2022). Note that the trend holds even when the architecture ultimately becomes locality-free (right-most). With the longest sequence length, it achieves the best accuracy.
Together, the two trends in Fig. 2 augment the observations from prior work that mainly focused on sufficiently large input sizes (Beyer et al., 2022; Hu et al., 2022), and give a more complete picture.
4.2 Case Study #2: Self-Supervised Learning
Next, we study the impact of removing locality with self-supervised pre-training and then fine-tuning for classification. In particular, we stick to ViT and choose MAE (He et al., 2022) for pre-training, thanks to its efficiency in computation and effectiveness for fine-tuning based evaluation protocols.
Setup.
We use CIFAR-100 (Krizhevsky, 2009) due to its inherent size of $\times32$. This allows us to fully explore the use of pixels as tokens on the original resolution. After pre-training, we test the classification performance on the same set, reporting top-1 (Acc@1) and top-5 (Acc@5) accuracy. For more details, please see Appendix A.
pre-train Acc@1 Acc@5
ViT-T/1 85.1 96.4
@@ 4.2 Case Study #2: Self-Supervised Learning
Table 4: Results for case study #2: self-supervised learning. We still use ViT (Dosovitskiy et al., 2021) as the model architecture, but explore MAE (He et al., 2022) as a pre-training technique for enhanced performance. The observation that locality-free variants (in bold) work well still holds with pre-training and the two model sizes we experimented.
(a) DiT-L/2 generations.
(b) DiT-L/1 generations.
Figure 3: Qualitative results for case study #3: image generation. These $\times256$ samples are generated from ImageNet-trained DiTs (Peebles & Xie, 2023). For direct comparisons, we fix random seeds and categories to prompt the model (none of the people classes from ImageNet are used), with the only difference that (a) uses locality-biased DiT-L/2, and (b) uses the locality-free variant (DiT-L/1). Overall, generations from locality-free models have fine features and detailed and reasonable, similar to locality-based models.
Results.
As shown in Tab. 4, the observation that locality-free variants of ViT work well still holds with MAE pre-training and the two model sizes we experimented.
4.3 Case Study #3: Image Generation
[…]
@@ 4.3 Case Study #3: Image Generation
The generation quality is measured by standard metrics: Fréchet Inception Distance (FID) (Heusel et al., 2017) with 50K samples, sFID (Nash et al., 2021), Inception Score (IS) (Salimans et al., 2016), and precision/recall (Kynkäänniemi et al., 2019), using reference batches from the original TensorFlow evaluation suite of Dhariwal & Nichol (2021).
@@
Sampled generations from DiT-L are shown in Fig. 3. The latent diffusion outputs are mapped back to the pixel space using the VQGAN decoder. A classifier-free guidance (Ho & Salimans, 2022) scale of 4.0 is used. For DiT/1, all generations are detailed and reasonable compared to the DiT/2 models with locality.
@@
DiT-XL/2 (Peebles & Xie, 2023), no guidance 10.67 - - - -
Table 5: Results for case study #3: image generation. We use references from the evaluation suite of Dhariwal & Nichol (2021), and report 5 metrics comparing locality-biased DiT-L/2 and locality-free DiT-L/1. Last row is from Peebles & Xie (2023), compared to which our baseline is significantly stronger (8.90 FID with DiT-L, vs. 10.67 with DiT-XL and longer training). Overall, our finding generalizes well to this new task with a different architecture and different input representations.
@@
Tab. 5 summarizes our observations. First, our baseline is strong: compared to the reference 10.67 FID (Peebles & Xie, 2023) with a larger model (DiT-XL/2) and longer training ($\sim$470 epochs), our DiT-L/2 achieves 8.90 without classifier-free guidance. Our main comparison (first two rows) uses a classifier-free guidance of 1.5. DiT-L/1, the locality-free variant, outperforms the baseline on all the 3 main metrics (FID, sFID and IS), and is on-par on precision/recall. With extended training of 1400 epochs, the gap is bigger: $2.68$ FID with DiT-L/1 vs. $2.89$ from DiT-L/2 (see Appendix C).
Our demonstration on image generation is an important extension of our finding. Compared to the case studies on discriminative benchmarks in Sec. 4.1 and Sec. 4.2, the task is more challenging; the model architecture is changed to be conditional; the input space is also changed from raw pixels to latent tokens. The fact that locality-free modification works out-of-the-box suggests the observation can be generalized across different tasks, architectural variants, and operating representations.
5 Locality Designs in ViT
Setup.
We use ViT-B for ImageNet supervised classification, adopting the exact same hyper-parameters, augmentations, and other training details from the scratch training recipe of He et al. (2022). Notably, images are crop-and-resized to $\times224$ and divided into $\times16$ patches.
Position embedding.
Similar to the investigation in Chen et al. (2021b), we choose from three candidates: sin-cos (Vaswani et al., 2017), learned, and none. The first option introduces locality into the model, while the other two do not. The results are summarized below:
PE sin-cos learned none
Acc@1 82.7 82.8 81.2
Our conclusion is similar to the one drawn by Chen et al. (2021b) for self-supervised representation evaluation: learnable position embeddings are on-par with fixed sin-cos ones. Interestingly, we observe only a minor drop in performance even if there is no position embedding at all – ‘none’ is only worse by 1.5% compared to sin-cos. Note that without position embedding, the classification model is fully permutation invariant w.r.t. patches, though not w.r.t. pixels.
@@ 5 Locality Designs in ViT
Figure 4: Pixel permutation for ViT. We swap pixels within a Hamming distance of $\delta$ and do this $T$ times (no distance constraint if $\delta=\inf$). Illustrated is an $\times8$ image divided into $\times2$ patches. Here we show permutation with $T=4$ pixel swaps (denoted by double-headed arrows).
@@
Next, we use learnable position embeddings and study patchification. To systematically reduce locality from patchification, our key insight is that neighboring pixels should no longer be tied in the same patch. To this end, we perform a pixel-wise permutation before diving the resulting sequence into separate tokens. Each token contains 256 pixels, same in number to pixels in a $\times16$ patch. The permutation is shared, staying the same for all the images including testing ones.
The permutation is performed in $T$ steps, each step will swap a pixel pair within a distance threshold $\delta\in[2,\inf]$ (22 means within the $\times2$ neighborhood, $\inf$ means any pixel pair can be swapped). We use hamming distance on the 2D image grid. $T$ and $\delta$ control how ‘corrputed’ an image is – larger $T$ or $\delta$ indicates more damage to the local neighborhood and thus more locality bias is taken away. Fig. 4 illustrates four such permutations.